VEGAS Team Meeting

March 16, 2012

Attendees

CASPER Team:
  • Dan Wertheimer
  • Mark Wagner
  • Glenn Jones
  • Andrew Siemion
  • Jayanth Chennamangalam

NRAO:
  • John Ford
  • Jason Ray
  • Patrick Brandt
  • Mark Whitehead
  • Randy McCullough
  • Marty Bloss
  • Anish Roshi
  • Paul Demorest

Meeting Notes:

No VEGAS teleconference on Monday, March 19

Observation Status: Decent weather for the galactic center observations in not expected in Green Bank for at least the next couple of weeks so observations cannot take place while the Berkeley/Caltech/WVU team is on site. It is possible to leave the current VEGAS setup in place for 2-3 more weeks at which time the borrowed roaches and ADCs must be returned. Green Bank agreed to supporting remote observations by the team with GB staff in place monitoring VEGAS during setup and execution of the observations. Since the KFPA wideband mode support in the IF Manager and Config Tool are not yet in place, setup will be managed by continuing to use 'set values' for broadband KFPA observations. Once the observations are complete or the time for borrowed ADCs and Roach boards has expired, GB will pack and return the items to their respective locations.

Testing Status: The overnight observations went pretty well and a set of bandpass and cross-polarization graphs were circulated to the team and the Green Bank science staff. There is a High frequency rolloff with a slope of ~6-10db. This rolloff may be cumulative ADC, IF, and cables which can be compensated for, but certainly should be investigated. Mark Wagner is working on a Mode 1 block diagram for the system documentations and also on a bug tracking system & version control, possibly using issue tracker capabilities on GitHub.

The new roach 'VEGAS-R9' is now in lab so the full bank of roaches is now available.

Other items:
  • The Advance Sig Ref provides a switching signal advanced in time. It is adjustable via pattern changes in state machine. This will be used to trigger items with an inherent latency (e.g. nutator). It is currently not connected outside and needs to be enabled via FPGA changes. This is a low priority task. The second 'sig ref' is for flagging internal data only, the LO frequency may be adjusted for these devices (??)
  • Once the loaned devices have been returned, four roaches will remain in the system in Green Bank
  • Mode 13 software changes - need to run different HPC accumulation program and adjust how to accept sub-band frequencies

With no observations in the near future, what else could we be doing?
  1. Measure the analog IF slope at input to VEGAS
  2. ADC tests National/!AsiaA
    • given AsiaA is looking like the best option for VEGAS purchase, it is best not focus on National ADC's 'bit glitch'
    • be aware that one AsiaA board has a shorted LSB. This may prevents doing crosstalk measurements unless we "short" other ADC's LSB. Green Bank techs will look for any shorted traces on the PCB. The other tests should be okay to run
  3. Data writing tests
  4. Switching signal generator troubleshooting (if not already solved this morning)
  5. K-band observations of some known pulsars
  6. Continuum observations

MartyBloss - 2012-03-16
Topic revision: r1 - 2012-03-16, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback